# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

### v2

@@ Appendix E Extend to Riemannian mix-max problems
which can be seen as a special case of Riemannian bilevel optimization problem (1) with $g(x,y)=-f(x,y)$.
If $f$ is geodesic strongly convex w.r.t. $y$ over ${\mathcal{M}}_{y}$, the Riemannian hypergradient (3) reduced to ${\mathcal{G}}F(x)={\mathcal{G}}_{x}f(x,y^{*}(x))$, and the approximate Riemannian hypergradient becomes $\widehat{{\mathcal{G}}}F(x,\hat{y})={\mathcal{G}}_{x}f(x,\hat{y})$. Then, the pseudocode of the adaptive method for solving Problem (6) is provided in Algorithm 4.
Algorithm 4 Riemannian bilevel solver for min-max optimization
1:Initia points $x_{0}\in{\mathcal{M}}_{x},y_{0}\in{\mathcal{M}}_{y}$, initial step sizes $a_{0}>0$, $b_{0}>0$, total iterations $T$, error tolerances $\epsilon_{y}=\frac{1}{T}$.
2:for $t=0,1,2,...,T-1$ do
3:     Set $p=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
@@
5:         $b_{k+1}^{2}=b_{k}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}$,
6:         $y_{t}^{k+1}={\rm Exp}_{y_{t}^{k}}(-\frac{1}{b_{k+1}},{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,
7:         Or $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}(-\frac{1}{b_{k+1}},{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$$\triangleright$ retraction mapping
8:         $p=p+1.$
@@
10:     $K_{t}=p$,
11:     $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})$,
12:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}$,
13:     $x_{t+1}={\rm Exp}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})$,
14:     Or $x_{t+1}={\rm Retr}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}}).$$\triangleright$ retraction mapping
@@
Before providing the convergence result of Algorithm 4, we claim the adopted assumptions required to converge as follows.
@@
    Assumptions 3.1 and 3.4 hold (Assumptions 3.5 and 3.6 hold when employing retraction);
@@
    Assumption 3.2 reduces to $f(x,y)$ is continuously differentiable and $-f(x,y)$ is geodesic strongly convex w.r.t. $y$;
@@
    Algorithm 3.3 reduces to $f(x,y)$ is $l_{f}$-Lipschitz continuous. The Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous.
We propose the following convergence result of Algorithm 4 without proof.
@@
where $C_{mm}$ is a constant relevant to the geodetic strong convexity, Lipschitzness, and curvature parameters.
Furthermore, the gradient complexity of Algorithm 4 is $G_{f}=O(1/\epsilon^{2})$.
@@ Appendix F Preliminary results for proofs
[78, Lemma 3.2] For any non-negative $\alpha_{1},...,\alpha_{T}$, and $\alpha_{1}\geq 1$, we have
@@
The proofs of Lemma F.2 can be derived from [24, Lemma 2.2], here we provide them for completeness.
@@
[…]
@@
$\leq$ $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{x}f(x_{2},y^{*}(x_{2}))\|_{x_{2}}$
$+\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{xy}^{2}g(x_{2},y^{*}(x_{2})){\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}\|_{x_{2}}\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\|{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}.$ (7)
@@
$\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))\|_{y^{*}(x_{2})}+\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
@@
$\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))-{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1}))\|_{y^{*}(x_{1})}+\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
@@
where the last inequality follows from the operator norm satisfies
@@
The remaining proof is similar to those of Lemma F.2 (2), thus, we omit it.
@@
By the terminated conditions of Algorithm 1 for the lower-level problem and the linear system, we have
@@
where the last inequality follows froms Assumptions 3.1, 3.2, 3.3, Lemmas F.3 and F.4. ∎
@@
$\leq$ $(L_{f}+\frac{\rho l_{f}}{\mu})d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}},$ (9)
@@
$\leq$ $\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})d(y_{t}^{K_{t}},y^{*}(x_{t})),$
@@
Combining it with (G.1) demonstrates that
@@
$\leq$ $\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\left(\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})d(y_{t}^{K_{t}},y^{*}(x_{t}))\right)$
@@ G.2.1 Improvement on objective function for one step update
$=$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x)\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
@@
Then, similar to [81, Lemma 8], we have the following upper bound for the step size $a_{t}$.
@@
[…]
@@
Removing the nonnegative term $-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}$, we have
@@
$C_{b}:=\max\{2\zeta L_{g}\frac{L_{g}}{\mu},b_{0}\},~{}~{}C_{c}:=\max\{L_{g},c_{0}\}.$ (18)
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})).$
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{2}{b_{k+1}}\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\rangle_{y_{t}^{k}}$
@@
$\leq$ $(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})),$
@@
$g(x_{t},y^{*}(x_{t}))\leq g(x_{t},y_{t}^{k})-\frac{1}{L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{L_{g}}{2}\|\frac{1}{L_{g}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}=g(x_{t},y_{t}^{k})-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}},$
@@
By the geodesic convex of $g$ w.r.t. $y$, we have
@@
If $k_{1}$ in Proposition G.1 does not exist, it holds that $b_{K_{t}}\leq C_{b}$. By [80, Lemma 2], we must have $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, since $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{k}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ hold for all $k<K^{t}$, we have
@@
where (a) follows from (2), (b) follows from the warm start of $y_{t}^{0}$, (c) follows from [7, Definition 10.1] and Lemma F.1, (d) uses Lemmas F.2 (1) and F.4, (e) follows from Lemma F.5 and $b_{k_{1}}\leq C_{b}$.
@@
where the second inequality follows from $b_{K}\geq C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$ and $1-x\leq e^{-x}$ for $10<x<$ 1, specifically, we have $C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$ and it holds that $10<\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}<$ 1, and the third inequality follows from (G.2.1).
@@
Denote $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right)$. Then, under (G.2.1), considering the monotonicity of $\frac{\mu}{b}-\zeta\frac{L_{g}^{2}}{b^{2}}$ w.r.t. $b$ when $b\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, it holds that $0<\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}\leq\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}$ as $\bar{b}\geq b_{K}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, we have
@@
If $n_{1}$ in Proposition G.1 does not exist, we haves $c_{N_{t}}\leq C_{c}$. Similar to $K_{t}$ of the lower-level problem, we have $N_{t}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$. If $N_{t}>\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$, by $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}\geq\epsilon_{v}$ and $c_{n}\leq C_{c}$ hold for all $n<N_{t}$, we have
@@
By the update mode of $v_{t}^{n_{1}}$ and the definition of $\hat{v}^{*}(x_{t},y_{t}^{K_{t}})=\operatorname*{arg\,min}_{v\in T_{y_{t}^{K_{t}}}{\mathcal{M}}_{y}}R(x_{t},y_{t}^{K_{t}},v)$, we have

### v3

@@ Appendix E Extension to Riemannian min-max Problems
which can be regarded as a special case of the Riemannian bilevel optimization problem (1) with g(x,y)=f(x,y)g(x,y)=-f(x,y).
[…]
Algorithm 4 AdaRHD for Riemannian Min-Max Optimization
1:Input: initial points $x_{0}\in{\mathcal{M}}_{x}$, $y_{0}\in{\mathcal{M}}_{y}$; initial step sizes $a_{0}>0$, $b_{0}>0$; total iterations $T$; tolerance $\epsilon_{y}=\frac{1}{T}$.
2:for $t=0,1,2,\ldots,T-1$ do
3:  Set $k=0$ and initialize $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$, otherwise $y_{t}^{0}=y_{0}$.
@@
5:   Update $b_{k+1}^{2}=b_{k}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}$.
6:   $y_{t}^{k+1}={\rm Exp}_{y_{t}^{k}}\!\left(-\frac{1}{b_{k+1}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\right)$,
7:         or $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}\!\left(-\frac{1}{b_{k+1}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\right).$ $\triangleright$ retraction step
8:   $k\leftarrow k+1$.
@@
10:  $K_{t}\leftarrow k$.
11:  Compute $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})$.
12:  Update $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}$.
13:  $x_{t+1}={\rm Exp}_{x_{t}}\!\left(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})\right)$,
14:         or $x_{t+1}={\rm Retr}_{x_{t}}\!\left(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})\right).$ $\triangleright$ retraction step
@@
Before establishing the convergence result of Algorithm 4, we present the required assumptions.
@@
    Assumptions 3.1 and 3.4 are satisfied (Assumptions 3.5 and 3.6 are required when retraction is used);
@@
    $f(x,y)$ is continuously differentiable, and f(x,y)-f(x,y) is geodesically strongly convex with respect to $y$;
@@
    $f(x,y)$ is $l_{f}$-Lipschitz continuous, and its Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous.
We now state the convergence guarantee of Algorithm 4.
@@
where $C_{mm}$ is a constant depending on the curvature, Lipschitz, and strong convexity parameters.
Moreover, the overall gradient complexity of Algorithm 4 is $\mathcal{O}(1/\epsilon^{2})$.
@@ Appendix F Preliminary results for proofs
[89, Lemma 3.2] For any non-negative $\alpha_{1},...,\alpha_{T}$, and $\alpha_{1}\geq 1$, we have
@@
The proofs of Lemma F.2 can be derived from [27, Lemma 2.2], here we provide them for completeness.
@@
$d(y^{*}(x_{1}),y^{*}(x_{2}))=\int_{0}^{1}\|c_{1}^{\prime}(t)\|_{c_{1}(t)}dt=\int_{0}^{1}\|{\mathrm{D}}y^{*}(c_{2}(t))[c_{2}^{\prime}(t)]\|_{c_{2}(t)}dt\leq\frac{L_{g}}{\mu}\int_{0}^{1}\|c_{2}^{\prime}(t)\|_{c_{2}(t)}dt=\frac{L_{g}}{\mu}d(x_{1},x_{2}),$
@@
$\leq$ $\left\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{x}f(x_{2},y^{*}(x_{2}))\right\|_{x_{2}}$
$+\left\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{xy}^{2}g(x_{2},y^{*}(x_{2})){\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}\right\|_{x_{2}}\left\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}\right\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\|{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}.$ (7)
@@
$\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}$
$\leq$ $\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}+\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
@@
$\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}$
$\leq$ $\left\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))-{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1}))\right\|_{y^{*}(x_{1})}+\left\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\right\|_{y^{*}(x_{2})}$
@@
where the last inequality follows from the operator norm satisfying
@@
The remaining proof is similar to that of Lemma F.2 (2), thus, we omit it.
@@
According to the termination conditions of Algorithm 1 for the lower-level problem and the linear system, we have
@@
where the last inequality follows from Assumptions 3.1, 3.2, 3.3, Lemmas F.3 and F.4. ∎
@@
$\leq$ $\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}},$ (9)
@@
$\leq$ $\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+\left(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}}\right)d(y_{t}^{K_{t}},y^{*}(x_{t})),$
@@
Combining this result with (G.1), we obtain
@@
$\leq$ $\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\left(\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+\left(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))\right)$
@@ G.2.1 Improvement on objective function for one step update
$=$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
@@
Then, similar to [92, Lemma 8], we have the following upper bound for the step size $a_{t}$.
@@
$\left\{\begin{aligned} a_{t}\leq&C_{a},~~&t\leq t_{1};\\ a_{t}\leq&C_{a}+2F_{0}+\frac{2t\epsilon_{y,v}}{a_{0}},~~&t>t_{1},\end{aligned}\right$.
@@
Removing the nonnegative term $\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}$, we have
@@
$C_{b}:=\max\left\{2\zeta L_{g}\frac{L_{g}}{\mu},b_{0}\right\},~~C_{c}:=\max\{L_{g},c_{0}\}.$ (18)
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq\left(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}}\right)d^{2}(y_{t}^{k},y^{*}(x_{t})).$
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))$ (19)
$\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{2}{b_{k+1}}\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\rangle_{y_{t}^{k}}$
@@
$\leq$ $\left(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}}\right)d^{2}(y_{t}^{k},y^{*}(x_{t})),$
@@
$g(x_{t},y^{*}(x_{t}))\leq g(x_{t},y_{t}^{k})-\frac{1}{L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{L_{g}}{2}\left\|\frac{1}{L_{g}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\right\|^{2}_{y_{t}^{k}}=g(x_{t},y_{t}^{k})-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}},$
@@
By the geodesic convexity of $g$ w.r.t. $y$, we have
@@
If $k_{1}$ in Proposition G.1 does not exist, it holds that $b_{K_{t}}\leq C_{b}$. By [91, Lemma 2], we must have $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, since $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{k}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ hold for all $k<K^{t}$, we have
@@
where (a) follows from (2), (b) follows from the warm start of $y_{t}^{0}$, (c) follows from [10, Definition 10.1] and Lemma F.1, (d) uses Lemmas F.2 (1) and F.4, (e) follows from Lemma F.5 and $b_{k_{1}}\leq C_{b}$.
@@
where the second inequality follows from $b_{K}\geq C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$ and $1-x\leq e^{-x}$ for $0<x<1$, specifically, when $b_{K}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, it holds that $0<\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}<1$, and the third inequality follows from (G.2.1).
@@
Denote $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right)$. We need to show that $\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}>0$. Under (G.2.1), considering the monotonicity of $\frac{\mu}{b}-\zeta\frac{L_{g}^{2}}{b^{2}}$ w.r.t. $b$ when $b\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, it holds that $0<\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}\leq\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}$ as $\bar{b}\geq b_{K}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$. Then, we have
@@
If $n_{1}$ in Proposition G.1 does not exist, we have $c_{N_{t}}\leq C_{c}$. Similar to $K_{t}$ of the lower-level problem, we have $N_{t}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$. If $N_{t}>\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$, by $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}\geq\epsilon_{v}$ and $c_{n}\leq C_{c}$ hold for all $n<N_{t}$, we have
@@
By the update rule of $v_{t}^{n_{1}}$ and the definition of $\hat{v}^{*}(x_{t},y_{t}^{K_{t}})=\operatorname*{arg\,min}_{v\in T_{y_{t}^{K_{t}}}{\mathcal{M}}_{y}}R(x_{t},y_{t}^{K_{t}},v)$, we have